OpenStudy (anonymous):

solve by the substitution method 3m+n=23 m-6n=14

OpenStudy (anonymous):

take the second equation and rearrange to get m=14+6n substitute for m in first equation ie: 3(14+6n)+n=23 42+19n=23 19n=-19 n=-1 substitute n=-1 in 3m+n=23 3m-1=23 3m=24 m=8 m=8 and n=-1